# A Reference-free Metric for Language-Queried Audio Source Separation using Contrastive Language-Audio Pretraining

### v1

@@ Abstract
Language-queried audio source separation (LASS) aims to separate an audio source guided by a text query, with the signal-to-distortion ratio (SDR)-based metrics being commonly used to objectively measure the quality of the separated audio. However, the SDR-based metrics require a reference signal, which is often difficult to obtain in real-world scenarios. In addition, with the SDR-based metrics, the content information of the text query is not considered effectively in LASS. This paper introduces a reference-free evaluation metric using a contrastive language-audio pretraining (CLAP) module, termed CLAPScore, which measures the semantic similarity between the separated audio and the text query. Unlike SDR, the proposed CLAPScore metric evaluates the quality of the separated audio based on the content information of the text query, without needing a reference signal. Experimental results show that the CLAPScore metric provides an effective evaluation of the semantic relevance of the separated audio to the text query, as compared to the SDR metric, offering an alternative for the performance evaluation of LASS systems.
@@ 1 Introduction
Following audio source separation literature [10, 11, 12], the signal-to-distortion ratio (SDR) based metrics, i.e., SDR [13], SDR improvement (SDRi) [14, 15], and scale-invariant SDR (SI-SDR) [16] have been used to measure the separation performance of LASS methods in [1]. All these metrics aim to quantify the quality of the separated audio signals. They measure how close the separated audio is to the original target audio, focusing on the reduction of distortion or errors introduced during the separation process [14].
@@
In this paper, we introduce a reference-free evaluation metric for language-queried audio source separation, which calculates the audio-text similarity score using the contrastive language-audio pretraining (CLAP) module [18], termed CLAPScore. Unlike the previous SDR-based metrics that require a reference audio to measure the separation performance, the proposed CLAPScore metric evaluates the semantic similarity between the separated audio and the text query without needing a reference audio. This makes CLAPScore metric particularly useful for real-world applications where a reference audio may not be available. Furthermore, similar to SDRi, the improvement in CLAPScore (CLAPScore-i) from the mixture to the separated audio can reflect the improvement from LASS methods. Moreover, the CLAPScore is also expanded to incorporate the reference audio while it is available, denoted as RefCLAPScore.
@@ 2.1 Definition of SDR-based Metrics
In widely used SDR-based metrics, SDR measures the ratio of the power of the desired signal to the power of the distortion introduced by the separation process [13]. SDRi is an improvement metric that measures the difference in SDR before and after applying an audio source separation algorithm [14, 15]. SI-SDR normalizes the signals to make the evaluation independent of their amplitude, which is more robust for varying scales [16, 19, 20]. The definition of SDR, SDRi and SI-SDR can be presented as follows:
@@ 2.2 Limitation of SDR-based Metrics
According to the above definition of SDR-based metrics, it can be found that, these metrics all depend on the reference audio $\mathbf{s}$ to measure the separation performance of the LASS methods. However, this requirement can be only met in a simulation scenario, where the reference audio and the noise are known to simulate the mixture audio. Due to the lack of the reference audio, these SDR-based metrics cannot be usable to measure the separation performance in the real-world scenario, as illustrated in Figure 1.
@@ 3.1 Definition of Proposed CLAPScore Metric
Specifically, the audio embedding of the estimated audio $\hat{\mathbf{s}}$ (i.e., the separated audio signal) and the text embedding of the text query are obtained with the CLAP module [^1] https://huggingface.co/spaces/Audio-AGI/AudioSep/blob/main/checkpoint/music_speech_audioset_epoch_15_esc_89.98.pt, as follows,
@@ 4.4 Evaluation with Different Mixing Strategies
Figure 4: Illustration of the proposed CLAPScore metric for mixtures from different mixing strategies.

### v2

@@ Abstract
Language-queried audio source separation (LASS) aims to separate an audio source guided by a text query, with the signal-to-distortion ratio (SDR)-based metrics being commonly used to objectively measure the quality of the separated audio. However, the SDR-based metrics require a reference signal, which is often difficult to obtain in real-world scenarios. In addition, with the SDR-based metrics, the content information of the text query is not considered effectively in LASS. This paper introduces a reference-free evaluation metric using a contrastive language-audio pretraining (CLAP) module, termed CLAPScore, which measures the semantic similarity between the separated audio and the text query. Unlike SDR, the proposed CLAPScore metric evaluates the quality of the separated audio based on the content information of the text query, without needing a reference signal. Experiments show that the CLAPScore provides an effective evaluation of the semantic relevance of the separated audio to the text query, as compared to the SDR metric, offering an alternative for the performance evaluation of LASS systems. The code for evaluation is publicly available [^1] GitHub: https://github.com/LittleFlyingSheep/CLAPScore˙for˙LASS.
@@ 1 Introduction
Following audio source separation literature [10, 11, 12], the signal-to-distortion ratio based metrics, i.e., SDR [13], SDR improvement (SDRi) [14, 15], and scale-invariant SDR (SI-SDR) [16] have been used to measure the separation performance of LASS methods in [1]. All these metrics aim to quantify the quality of the separated audio signals. They measure how close the separated audio is to the original target audio, focusing on the reduction of distortion or errors introduced during the separation process [14].
@@
In this paper, we introduce a reference-free evaluation metric for LASS, which calculates the audio-text similarity score using the contrastive language-audio pretraining (CLAP) module [18], termed CLAPScore. Unlike the previous SDR-based metrics that require a reference audio to measure the separation performance, the proposed CLAPScore metric evaluates the semantic similarity between the separated audio and the text query without needing a reference audio. This makes CLAPScore metric particularly useful for real-world applications where a reference audio may not be available. Furthermore, similar to SDRi, the improvement in CLAPScore (CLAPScore-i) from the mixture to the separated audio can reflect the improvement from LASS methods. Moreover, the CLAPScore is also expanded to incorporate the reference audio while it is available, denoted as RefCLAPScore.
@@ 2.1 Definition of SDR-Based Metrics
In widely used SDR-based metrics, SDR measures the ratio of the power of the desired signal to the power of the distortion introduced by the separation process [13]. SDRi is an improvement metric that measures the difference in SDR before and after applying an audio source separation algorithm [14, 15]. SI-SDR normalizes the audio signals to make the evaluation independent of their amplitude, which is more robust for varying scales [16, 19, 20]. The definition of SDR, SDRi and SI-SDR can be presented as follows:
@@ 2.2 Limitation of SDR-Based Metrics
According to the above definition of SDR-based metrics, it can be found that, these metrics all depend on the reference audio signal $\mathbf{s}$ to measure the separation performance of the LASS methods. However, this requirement can be only met in a simulation scenario, where the reference audio and the noise are known to simulate the mixture audio. Due to the lack of the reference audio, these SDR-based metrics cannot be usable to measure the LASS performance in the real-world scenario, as illustrated in Figure 1.
@@ 3.1 Definition of Proposed CLAPScore Metric
Specifically, the audio embedding of the estimated audio $\hat{\mathbf{s}}$ (i.e., the separated audio signal) and the text embedding of the text query are obtained with the CLAP module [^2] https://huggingface.co/spaces/Audio-AGI/AudioSep/blob/main/checkpoint/music˙speech˙audioset˙epoch˙15˙esc˙89.98.pt, as follows,
@@ 4.4 Evaluation with Different Mixing Strategies
Figure 4: Illustration of the proposed CLAPScore metric for the mixtures from different mixing strategies.